# ChemGen: Code Generation for Multispecies Chemical Kinetics in Physics Simulations

### v1

@@ 2.2.3 Fall-off reactions
The same optimization for efficiencies can be made as in the third-body prefactor. Various falloff reactions exist where the only difference is in the function $F$, known as the broadening term, which we outline below, the simplest being $F=1$ from Lindemann [31].
@@
The origin of these fits is given in [32, 33, 34, 35].
@@
In the SRI [36] broadening approach, $F$ is represented by
@@
2.2.4 Pressure-dependent Arrhenius rate expressions
The previous fall-off reaction rates seek to fit between a low-pressure reaction rate and a high-pressure reaction rate. However, some rates have several pressure dependent zones that cannot be accurately represented by a single or two Arrhenius expressions. The pressure-dependent Arrhenius rate expression, known as the P-log reaction rate, was introduced to utilize pressure-dependent reaction rates by logarithmically interpolating between Arrhenius rate expressions at various pressures [37]. Given two Arrhenius reaction rates at two specific pressures,
@@ 2.2.4 Pressure-dependent Arrhenius rate expressions
An additional complexity is that multiple rate expressions may be given at the same pressure, in which case the rate used in the interpolation formula is the sum of all the rates given at that pressure. For pressures outside the given range, the rate expression at the nearest bounded pressure is used. In ChemGen, the $\ln p_{i}$ terms are calculated during code generation.
To further elaborate on this rate, consider, for example, the 81st reaction in FFCM2 [20],
@@ 3 Time integration strategies
The stability of the CFD time integration may be compromised by explicitly applying the source term of Equation (2). Different approaches can be utilized by time-splitting the evolution of the chemical state and retain the stability of the fluid dynamic algorithms [38, 39, 13, 15]. These approaches typically do not involve the flux from either convection or diffusion. By removing any spatial fluxes, the conservation Equation (1) reduces to an ordinary differential equation,
@@ 3.1 Explicit Runge-Kutta time integration method
ChemGen provides a Runge-Kutta 4th-order method (RK4) for integrating the chemical state from $y^{n}$ to $y^{n+1}$ over a time interval of $\Delta t$ from Equation (48). This explicit time integration scheme approximates the solution by evaluating the source term at four intermediate stages within each time step [40]. Given state $y_{n}$ at time $t_{n}$, the RK4 method updates to the next state, $y^{n+1}$ at time $t^{n+1}$. The algorithm is
@@ 3.2 Implicit time integration and Jacobians
$\mathcal{G}\left(y_{k}^{n+1}\right)\left(y_{k+1}^{n+1}-y_{k}^{n+1}\right)=-f(y_{k}^{n+1}),\quad f(y_{k}^{n+1})=\frac{y^{n+1}-y^{n}}{\Delta t}-S\left(y^{n+1}\right),$ (51)
@@
where $\mathcal{D}_{ij}$ depends on the time integration scheme ($\mathcal{D}_{ij}=\delta_{ij}/\Delta t$ for the backward Euler method) and $\mathcal{J}_{ij}$ is the Jacobian of the chemical source term with $y_{j}$ as the $j$th element of $y$ and $S_{i}$ is the $i$th element of of $S$.
@@
$\mathcal{J}_{ij}=\begin{cases}\frac{\partial\omega_{T}}{\partial T}&j=1\\ \frac{\partial\omega_{T}}{\partial C_{j-1}}&i=1;\quad j>1\\ \frac{\partial\omega_{i-1}}{\partial T}&i>1;\quad j=1\\ \frac{\partial\omega_{i-1}}{\partial C_{j-1}}&i>1;j>1\end{cases},$ (55)
@@ 3.3 Additional time integrators
Other time integration strategies not yet implemented in ChemGen can be added using the same framework as the existing methods. These include, but are not limited to, multi-step implicit methods such as backward differentiation formulas, which utilize information from previous time steps, and higher-order Rosenbrock methods [41], which require additional stages. Integration methods such as DGODE [15] and CHEMEQ2 [42] may be more challenging to implement. DGODE is based on finite element numerics, while CHEMEQ2 requires the source term to be separated into distinct production and consumption components, introducing additional complexities not currently addressed by ChemGen.
@@ 3.4 Linear solvers
It is well-established that employing a direct solver requires $\mathcal{O}\left(n^{3}\right)$ operations for an $n\times n$ square matrix [43]. Although ChemGen supplies a direct solver, its default linear solver is a stand alone function that uses the Generalized Minimal Residual (GMRES) method in order to mitigate the $\mathcal{O}(n^{3})$ cost [44]. GMRES is an iterative Krylov subspace method that builds an orthonormal basis of the Krylov subspace and minimizes the residual over this subspace at each iteration.
@@
ChemGen provides two standard options: Jacobi preconditioning and Gauss-Seidel preconditioning. In the former, $M$ simply consists of the diagonal of $A$, which can be efficiently inverted by computing the reciprocal of each diagonal entry. In the latter, $M$ consists of the lower triangular portion of $A$ (including the diagonal), which can be easily applied with forward substitution. For more advanced preconditioning techniques used in chemical kinetics, see, for example, [45].
ChemGen is not intended to serve as a comprehensive linear-solver library. For more advanced or scalable linear solver capabilities, ChemGen can be interfaced with external libraries such as Eigen [46] or PETSc [47], which offer efficient preconditioned Krylov solvers and direct sparse solvers for large systems. Additionally, while ChemGen includes several implicit time integration strategies (e.g., backward Euler, SDIRK methods), offloading the integration of stiff ODE systems to established libraries like CVODES [48] is also a reasonable approach. Such packages allow user-supplied source terms and Jacobians that can then be used in time integration strategies.
@@ 4.1 The concept of decorators
to be declared as a const in a struct or class. In addition, the scalar float could also be double, and the user may wish to pass the function parameters by const reference, i.e., float a $\rightarrow$ const float& a. A function annotation may also be desirable, such as a preprocessor macro provided the Kokkos library [49]. This can turn the above example to
@@ 4.2 Derivatives
The mechanical differentiation approach is both efficient and more readable than symbolic differentiation using tools such as SymPy [50]. While symbolic differentiation can be valuable, the mechanical approach is better suited for organizing and maintaining ChemGen’s code generation process. Regardless, since the functions are available in C++ , packages such as autodiff [51] can be implemented to use a wider variety of differentiable functionality. For a mathematical overview of the derivatives needed for chemically reacting flow, see [24].
@@ 5 Results and demonstration cases
This section presents standalone test cases to confirm that the generated source code is behaving as intended. We then embed the generated code in detonationEulerFoam [52] based on the OpenFOAM package [53] to demonstrate its performance in a CFD simulation of a moving detonation wave.
5.1 Source term accuracy analysis
@@
where are random values $r_{T},r_{p},r_{1},\dots,r_{n_{s}}$ are generated from a uniform distribution, $\sim\mathcal{U}(0,1)$. $X_{i}$ is the mole fraction of the $i$th species, defined as
@@
Ó Connaire 10 40 [54]
Burke 13 27 [55]
GRI-Mech 3.0 53 325 [56]
UCSD 63 268 [57]
FFCM-2 96 1054 [20]
@@
where $S_{\mathrm{ct},i}$ denotes the Cantera-predicted source term for the $i$th specie and $S_{\mathrm{cg},i}$ denotes the corresponding ChemGen-predicted source term. If the magnitude of the Cantera source term is less than $10^{-10}$, the error is considered to be zero.
@@ 5.2 Explicit time integration
We use a hydrogen submodel extracted from FFCM-2 [20], which is created by removing carbon containing species, electronically excited states, such as $OH^{*}$, and inert species $Ar$ and $He$. The resulting submodel consists of 9 species and 25 reactions. We calculate the concentrations, $C_{i}$, in the same manner as in Section (5.1). We employ a time-step size of $\Delta t=1\times 10^{-8}$ seconds. With larger time-step sizes, the RK4 time integrator becomes unstable. The resulting temperature and species concentration profiles are shown in Figure (3) , where the ChemGen results are represented by dashed black lines and Cantera results are represented by solid colored lines. We produce the Cantera results using the ReactorNet class.
Figure 3: Temperature (left) and species (right) evolution for a constant volume reactor using an explicit RK4 scheme and initial conditions given in Equation (63). These results are produced using a hydrogen submodel extracted from FFCM2 [20]. The solid lines represent the Cantera solution, and the dashed black lines represent the ChemGen solution.
@@ 5.3 Jacobian verification
$\hat{\mathcal{J}_{ij}}$ $=$ $\begin{cases}\frac{\omega_{T}(C,T+\delta T)-\omega_{T}(C,T-\delta T)}{2\delta T}&j=1\\ \frac{\omega_{T}(C_{j-1}^{+},T)-\omega_{T}(C_{j-1}^{-},T)}{2\delta C}&i=1;\quad j>1\\ \frac{\omega_{i-1}(C,T+\delta T)-\omega_{i-1}(C,T-\delta T)}{2\delta T}&i>1;\quad j=1\\ \frac{\omega_{i-1}\left(C_{j-1}^{+},T\right)-\omega_{i-1}\left(C_{j-1}^{-},T\right)}{2\delta C}&i>1;j>1,\end{cases}$ (67)
@@
for four refinement levels using consecutively halved values of $\delta C$ and $\delta T$. The initial $\delta C$ is chosen as a quarter of the smallest non-zero concentration, and the initial temperature perturbation is chosen as $\delta T=100$ K. Figure (4) shows the order of convergence for a single random chemical state determined using the UCSD chemical model [57]. The refinement results agree well with the expected second-order accuracy. We repeat this refinement study across all previously mentioned chemical models and observed consistent order of accuracy for both Jacobians (not shown for brevity). In some cases, the initial $\delta C$ is too small, leading to errors approaching machine precision and preventing further reduction in the $L_{2}$-norm. By adjusting the initial $\delta C$, the expected order of accuracy was recovered.
@@
$\mathcal{\hat{J}}_{ij,1}$ $=$ $\begin{cases}\frac{\omega_{T}(C,T+\delta T)-\omega_{T}(C,T)}{\delta T}&j=1\\ \frac{\omega_{T}(C_{j-1}^{+},T)-\omega_{T}(C,T)}{\delta C}&i=1;\quad j>1\\ \frac{\omega_{i-1}(C,T+\delta T)-\omega_{i-1}(C,T)}{\delta T}&i>1;\quad j=1\\ \frac{\omega_{i-1}\left(C_{j-1}^{+},T\right)-\omega_{i-1}\left(C,T\right)}{\delta C}&i>1;j>1.\end{cases}$ (69)
The results are summarized in Table 4 for a variety of chemical models ranging in size and complexity. Despite using the fastest discrete Jacobian optimization, the analytical Jacobians are consistently faster across all cases. As the chemical mechanism increases in complexity, the reduction in cost of computing the analytical Jacobian relative to the finite different approximation becomes greater.
@@
Finally, Niemeyer et al. [24] reported error metrics that can be used to verify Jacobian accuracy. These metrics are utilized in C.2 to provide additional confidence in the correctness of ChemGen’s Jacobian computations.
@@ 5.4 Implicit time integration
This section presents results obtained with the implicit time integrators provided by ChemGen, motivated by the confirmed accuracy of the source term from Section 5.1, successful explicit time integration from Section 5.2, and confidence in the computed source term Jacobian from Section 5.3. We utilize the FFCM-2 chemical model [20] to integrate a chemical state with initial conditions
@@ 5.5 ChemGen in a detonation simulation
To demonstrate the usage of ChemGen in an existing computational framework, we target the chemistry evaluation in detonationFoam [52] which is based on the OpenFOAM package [53]. In particular, we focus on the detonationEulerFoam solver, which does not include viscous effects. The solver treats each individual species conservation equation via a segregated approach by solving
@@
where $\nabla_{\mathrel{FVM}}\cdot\mathcal{F}\left(\rho Y_{i}\right)$ denotes the finite volume discretization of the species flux and $\mathcal{R}_{i}$ is an applied source term of the $i$th specie. For the complete set of conservation equations and additional details on the computational strategy, see [52]. Here the source term is calculated as
@@
In detonationFoam [52], the seulex ODE solver [40] supplied by OpenFOAM is used to solve for $C_{i}^{n+1}$. The seulex solver utilizes an extrapolation algorithm based on the implicit Euler method with step size control and order selection. Due to the small time step size $\Delta t=1\times 10^{-9}$ s we replace the seulex solver with ChemGen’s YASS integrator combined with Jacobi-preconditioned GMRES.
We consider a two-dimensional detonation test that has been simulated in several previous studies (see Oran et al. [58], Houim and Kuo [39], Lv and Ihme [13], Johnson and Kercher [15], and Ching et al. [17]). The domain here, which differs slightly from those in the aforementioned works, is $0.5$ m long ($x_{1}$-direction) with a channel height of $0.06$ m ($x_{2}$-direction). The initial mesh is a uniform quadrilateral mesh with $2000$ cells in the $x_{1}$-direction and $240$ cells in $x_{2}$-direction. We use the AMR default settings from [52], resulting in three possible refinements, with a minimum mesh size of $31\times 10^{-6}$ m. We use the slip wall condition for all boundaries. We again utilize the hydrogen submodel extracted from FFCM-2 [20]. The initial conditions are based on a steady detonation solution from the Shock & Detonation Toolbox [59], where the pre-shock state is given by
@@
In addition, we perturb the detonation front such that it varies in the $x_{2}$-direction. In Sun et al. [52] and in this work a sinusoidal profile is specified to perturb the front location, giving rise to instabilities that then transition to the expected solution.

### v3

@@ 2.2.3 Pressure fall-off effect
The same optimization for efficiencies can be made as in the third-body prefactor. Various falloff reactions exist where the only difference is in the function $F$, known as the broadening term, which we outline below, the simplest being $F=1$ from Lindemann [Lin22].
@@
The origin of the empirical parameters is given in [Tro77, Tro79, Gil83, Tro87].
@@
In the SRI [Ste89] broadening approach, $F$ is represented by
@@
2.2.4 Other forms of pressure-dependent reactions
Some reaction rates have several pressure-dependent expressions that cannot be accurately represented by single Arrhenius expressions. Known as the P-log reaction rate, logarithmically interpolating between Arrhenius rate expressions at various pressures [Gou11_b] is required. Specifically, given two Arrhenius rates at two specific pressures,
@@ 2.2.4 Other forms of pressure-dependent reactions
An additional complexity is that multiple rate expressions may be given at the same pressure because of non-Arrhenius behavior, in which case the rate used in the interpolation formula is the sum of all the rates given at that pressure. For pressures outside the given range, the rate expression at the nearest bounded pressure is used. In ChemGen, the $\ln p_{i}$ terms are calculated during code generation.
To further elaborate on this rate, consider, for example, the 81st reaction in FFCM2 [Zha23],
@@ 3 Time integration strategies
The stability of the CFD time integration may be compromised by explicitly applying the source term of Equation (2). Different approaches can be utilized by time-splitting the evolution of the chemical state and retain the stability of the fluid dynamic algorithms [Str68, Hou11, Lv15, Joh20]. These approaches typically do not involve the flux from either convection or diffusion. By removing any spatial fluxes, the conservation Equation (1) reduces to an ordinary differential equation,
@@ 3.1 Explicit Runge-Kutta time integration method
ChemGen provides a Runge-Kutta 4th-order method (RK4) for integrating the chemical state from $y^{n}$ to $y^{n+1}$ over a time interval of $\Delta t$ from Equation (49). This explicit time integration scheme approximates the solution by evaluating the source term at four intermediate stages within each time step [Hai96]. Given state $y_{n}$ at time $t_{n}$, the RK4 method updates to the next state, $y^{n+1}$ at time $t^{n+1}$. The algorithm is
@@ 3.2 Implicit time integration and Jacobians
$\mathcal{G}\left(y_{k}^{n+1}\right)\left(y_{k+1}^{n+1}-y_{k}^{n+1}\right)=-f(y_{k}^{n+1}),\quad f(y_{k}^{n+1})=\frac{y^{n+1}_{k}-y^{n}}{\Delta t}-S\left(y_{k}^{n+1}\right),$ (52)
@@
where $\mathcal{D}_{ij}$ depends on the time integration scheme ($\mathcal{D}_{ij}=\delta_{ij}/\Delta t$ for the backward Euler method) and $\mathcal{J}_{ij}$ is the Jacobian of the chemical source term with $y_{j}$ as the $j^{th}$ element of $y$ and $S_{i}$ is the $i^{th}$ element of of $S$.
@@
$\mathcal{J}_{ij}=\begin{cases}\frac{\partial\omega_{T}}{\partial T}&i=1;\quad j=1\\ \frac{\partial\omega_{T}}{\partial C_{j-1}}&i=1;\quad j>1\\ \frac{\partial\omega_{i-1}}{\partial T}&i>1;\quad j=1\\ \frac{\partial\omega_{i-1}}{\partial C_{j-1}}&i>1;j>1\end{cases},$ (56)
@@ 3.3 Additional time integrators
Other time integration strategies not yet implemented in ChemGen can be added using the same framework as the existing methods. These include, but are not limited to, multi-step implicit methods such as backward differentiation formulas, which utilize information from previous time steps, and higher-order Rosenbrock methods [Zha14], which require additional stages. Integration methods such as DGODE [Joh20] and CHEMEQ2 [Mot01] may be more challenging to implement. DGODE is based on finite element numerics, while CHEMEQ2 requires the source term to be separated into distinct production and consumption components, introducing additional complexities not currently addressed by ChemGen.
@@ 3.4 Linear solvers
It is well-established that employing a direct solver requires $\mathcal{O}\left(n^{3}\right)$ operations for an $n\times n$ square matrix [Gol13]. Although ChemGen supplies a direct solver, its default linear solver is a stand alone function that uses the Generalized Minimal Residual (GMRES) method in order to mitigate the $\mathcal{O}(n^{3})$ cost [Saa86]. GMRES is an iterative Krylov subspace method that builds an orthonormal basis of the Krylov subspace and minimizes the residual over this subspace at each iteration.
@@
ChemGen provides two standard options: Jacobi preconditioning and Gauss-Seidel preconditioning. In the former, $M$ simply consists of the diagonal of $A$, which can be efficiently inverted by computing the reciprocal of each diagonal entry. In the latter, $M$ consists of the lower triangular portion of $A$ (including the diagonal), which can be easily applied with forward substitution. For more advanced preconditioning techniques used in chemical kinetics, see, for example, [Wal23].
ChemGen is not intended to serve as a comprehensive linear-solver library. For more advanced or scalable linear solver capabilities, ChemGen can be interfaced with external libraries such as Eigen [Gue10] or PETSc [Bal97], which offer efficient preconditioned Krylov solvers and direct sparse solvers for large systems. Additionally, while ChemGen includes several implicit time integration strategies (e.g., backward Euler, SDIRK methods), offloading the integration of stiff ODE systems to established libraries like CVODES [Ser05] is also a reasonable approach. Such packages allow user-supplied source terms and Jacobians that can then be used in time integration strategies.
@@ 4.1 The concept of decorators
to be declared as a const in a struct or class. In addition, the scalar float could also be double, and the user may wish to pass the function parameters by const reference, i.e., float a $\rightarrow$ const float& a. A function annotation may also be desirable, such as a preprocessor macro provided the Kokkos library [Tro22]. This can turn the above example to
@@ 4.2 Derivatives
The mechanical differentiation approach is both efficient and more readable than symbolic differentiation using tools such as SymPy [sympy]. While symbolic differentiation can be valuable, the mechanical approach is better suited for organizing and maintaining ChemGen’s code generation process. Regardless, since the functions are available in C++ , packages such as autodiff [Lea18] can be implemented to use a wider variety of differentiable functionality. For a mathematical overview of the derivatives needed for chemically reacting flow, see [Nie17].
@@ 5 Results and demonstration cases
This section presents standalone test cases to confirm that the generated source code is behaving as intended. We then embed the generated code in detonationEulerFoam [Sun23] based on the OpenFOAM package [Ope25] to demonstrate its performance in a CFD simulation of a moving detonation wave.
5.1 Source term accuracy
@@
where are random values $r_{T},r_{p},r_{1},\dots,r_{n_{s}}$ are generated from a uniform distribution, $\sim\mathcal{U}(0,1)$. $X_{i}$ is the mole fraction of the $i^{th}$ species, defined as
@@
Ó Connaire 10 40 [Oco04]
Burke 13 27 [Bur12]
GRI-Mech 3.0 53 325 [Smi00]
UCSD 63 268 [SanDiegoMech]
FFCM-2 96 1054 [Zha23]
@@
where $S_{\mathrm{ct},i}$ denotes the Cantera-predicted source term for the $i^{th}$ specie and $S_{\mathrm{cg},i}$ denotes the corresponding ChemGen-predicted source term. If the magnitude of the Cantera source term is less than $10^{-10}$, the error is considered to be zero.
@@ 5.2 Explicit time integration
We use a hydrogen submodel extracted from FFCM-2 [Zha23], which is created by removing carbon containing species, electronically excited states, such as $OH^{*}$, and inert species $Ar$ and $He$. The resulting submodel consists of 9 species and 25 reactions. We calculate the concentrations, $C_{i}$, in the same manner as in Section (5.1). We employ a time-step size of $\Delta t=1\times 10^{-8}$ seconds. With larger time-step sizes, the RK4 time integrator becomes unstable. The resulting temperature and species concentration profiles are shown in Figure (3) , where the ChemGen results are represented by dashed black lines and Cantera results are represented by solid colored lines. We produce the Cantera results using the ReactorNet class.
Figure 3: Temperature (left) and species (right) evolution for a constant volume reactor using an explicit RK4 scheme and initial conditions given in Equation (64). These results are produced using a hydrogen submodel extracted from FFCM2 [Zha23]. The solid lines represent the Cantera solution, and the dashed black lines represent the ChemGen solution.
@@ 5.3 Jacobian verification
$\hat{\mathcal{J}_{ij}}$ $=$ $\begin{cases}\frac{\omega_{T}(C,T+\delta T)-\omega_{T}(C,T-\delta T)}{2\delta T}&i=1;\quad j=1\\ \frac{\omega_{T}(C_{j-1}^{+},T)-\omega_{T}(C_{j-1}^{-},T)}{2\delta C}&i=1;\quad j>1\\ \frac{\omega_{i-1}(C,T+\delta T)-\omega_{i-1}(C,T-\delta T)}{2\delta T}&i>1;\quad j=1\\ \frac{\omega_{i-1}\left(C_{j-1}^{+},T\right)-\omega_{i-1}\left(C_{j-1}^{-},T\right)}{2\delta C}&i>1;\quad j>1,\end{cases}$ (68)
@@
for four refinement levels using consecutively halved values of $\delta C$ and $\delta T$. The initial $\delta C$ is chosen as a quarter of the smallest non-zero concentration, and the initial temperature perturbation is chosen as $\delta T=100$ K. Figure (4) shows the order of convergence for a single random chemical state determined using the UCSD chemical model [SanDiegoMech]. The refinement results agree well with the expected second-order accuracy. We repeat this refinement study across all previously mentioned chemical models and observed consistent order of accuracy for both Jacobians (not shown for brevity). In some cases, the initial $\delta C$ is too small, leading to errors approaching machine precision and preventing further reduction in the $L_{2}$-norm. By adjusting the initial $\delta C$, the expected order of accuracy was recovered.
@@
$\mathcal{\hat{J}}_{ij,1}$ $=$ $\begin{cases}\frac{\omega_{T}(C,T+\delta T)-\omega_{T}(C,T)}{\delta T}&i=1;\quad j=1\\ \frac{\omega_{T}(C_{j-1}^{+},T)-\omega_{T}(C,T)}{\delta C}&i=1;\quad j>1\\ \frac{\omega_{i-1}(C,T+\delta T)-\omega_{i-1}(C,T)}{\delta T}&i>1;\quad j=1\\ \frac{\omega_{i-1}\left(C_{j-1}^{+},T\right)-\omega_{i-1}\left(C,T\right)}{\delta C}&i>1;j>1.\end{cases}$ (70)
The results are summarized in Table 4 for a variety of chemical models ranging in size and complexity. Despite using the fastest discrete Jacobian optimization, the analytical Jacobians are consistently faster across all cases. As the chemical reaction model increases in complexity, the reduction in cost of computing the analytical Jacobian relative to the finite different approximation becomes greater.
@@
Finally, Niemeyer et al. [Nie17] reported error metrics that can be used to verify Jacobian accuracy. These metrics are utilized in C.2 to provide additional confidence in the correctness of ChemGen’s Jacobian computations.
@@ 5.4 Implicit time integration
This section presents results obtained with the implicit time integrators provided by ChemGen, motivated by the confirmed accuracy of the source term from Section 5.1, successful explicit time integration from Section 5.2, and confidence in the computed source term Jacobian from Section 5.3. We utilize the FFCM-2 chemical model [Zha23] to integrate a chemical state with initial conditions
@@ 5.5 ChemGen in a detonation simulation
To demonstrate the usage of ChemGen in an existing computational framework, we target the chemistry evaluation in detonationFoam [Sun23] which is based on the OpenFOAM package [Ope25]. In particular, we focus on the detonationEulerFoam solver, which does not include viscous effects. The solver treats each individual species conservation equation via a segregated approach by solving
@@
where $\nabla_{\mathrel{FVM}}\cdot\mathcal{F}\left(\rho Y_{i}\right)$ denotes the finite volume discretization of the species flux and $\mathcal{R}_{i}$ is an applied source term of the $i^{th}$ specie. For the complete set of conservation equations and additional details on the computational strategy, see [Sun23]. Here the source term is calculated as
@@
In detonationFoam [Sun23], the seulex ODE solver [Hai96] supplied by OpenFOAM is used to solve for $C_{i}^{n+1}$. The seulex solver utilizes an extrapolation algorithm based on the implicit Euler method with step size control and order selection. Due to the small time step size $\Delta t=1\times 10^{-9}$ s we replace the seulex solver with ChemGen’s YASS integrator combined with Jacobi-preconditioned GMRES.
We consider a two-dimensional detonation test that has been simulated in several previous studies (see Oran et al. [Ora98], Houim and Kuo [Hou11], Lv and Ihme [Lv15], Johnson and Kercher [Joh20], and Ching et al. [Chi22_2]). The domain here, which differs slightly from those in the aforementioned works, is $0.5$ m long ($x_{1}$-direction) with a channel height of $0.06$ m ($x_{2}$-direction). The initial mesh is a uniform quadrilateral mesh with $2000$ cells in the $x_{1}$-direction and $240$ cells in $x_{2}$-direction. We use the AMR default settings from [Sun23], resulting in three possible refinements, with a minimum mesh size of $31\times 10^{-6}$ m. We use the slip wall condition for all boundaries. We again utilize the hydrogen submodel extracted from FFCM-2 [Zha23]. The initial conditions are based on a steady detonation solution from the Shock & Detonation Toolbox [sdtoolbox], where the pre-shock state is given by
@@
In addition, we perturb the detonation front such that it varies in the $x_{2}$-direction. In Sun et al. [Sun23] and in this work a sinusoidal profile is specified to perturb the front location, giving rise to instabilities that then transition to the expected solution.